OpenStudy (anonymous):

Helppppppp Write equations of the line through the given point a) parallel and b) normal to the given linen a.(5,-3), x+y=4 b.(-6,2), 5x+2y=7 c.(-3,-4), y=-2

OpenStudy (tkhunny):

Slope-Intercept form is normally helpful for problems requiring the slope. Careful with c.

OpenStudy (anonymous):

a.(5,-3), x+y=4 (parallel lines have same slpoe) Since line is passing through (5,-3) and is parallel to x+y=4 i.e. y= -x+4 comparing y= -x+4 by y = mx +c we find slope m= -1 eq of line in slope point form is given as y-y1=m(x-x1) y-(-3)=-1(x-5) y+3=-x+5 y= -x+5-3 y=-x+2 x+y=2 is the required eq of the line

OpenStudy (anonymous):

b.(-6,2), 5x+2y=7 Since line is passing through (-6,2) and is normal to the line 5x+2y=7 i.e. 2y = -5x+7 y=( -5/2)x+7/2 therefore m=-5/2 slope of req line = 2/5 now proceed in the same way as above

OpenStudy (anonymous):

@Gabysolis49